# MicroDreamer: Zero-shot 3D Generation in $\sim$20 Seconds by Score-based Iterative Reconstruction

### v2

@@ B.1 SIR enabling optimization in the pixel space
B.2 Faliure of applying SDS in pixel space.
We highlight that SDS optimizing in pixel space can not be effective even if it has the data predicting form by reparameterized as presented in Eq. (6). Fig. 9 shows the poor quality for single-step prediction from the diffusion model using Eq. (2). Therefore, if we attempt to employ SDS for optimization in the pixel space, i.e. optimizing 3D objects with 2D images like those generated in Fig. 9, the resulting 3D content would be of subpar quality, as the results presented in Fig. 10.
Figure 9: Results of single-step prediction. For each pair of images, the left is the input image and the right is the output of single-step prediction. The poor quality of the output makes it unsuitable for 3D optimization.
@@
Figure 10: Results of applying SDS in pixel space. SDS optimizing in pixel space can not be effective.
Appendix C Time schedule of $t_{1}$
The hyperparameter $t_{1}$ regulates the balance between inversion and noise-adding, which varies with different models. In our earlier exploration, we adopted $t_{1}=t^{2}_{2}/T$ for SIR based on MVDream [42] and ImageDream [44] on NeRF, and $t_{1}=0.6t_{2}$ for SIR based on Stable Zero123 [43, 6] on both NeRF and 3DGS. Notably, SIR based on ImageDream on 3DGS opted for $t_{1}=t_{2}$, i.e., adding noise for the forward process. As an advanced multi-view diffusion model, ImageDream can generate images with enhanced 3D consistency, obviating the need for inversion to maintain consistency. Fig. 11 shows some results of SIR on ImageDream on 3DGS using $t_{1}=t_{2}$ schedule.
Figure 11: Results of SIR based on ImageDream on 3DGS. These results images are rendered from 3DGS with no refinement stage.
Appendix D More implementation details
Codebase.
For the implementation of SIR and SDS on NeRF, we choose the popularly used framework threestudio [56]. For the implementation of SIR on 3DGS, we follow the framework from DreamGaussian [9]. The checkpoint we chose for CLIP computation is laion/CLIP-ViT-bigG-14-laion2B-39B-b160k [45, 77, 78]. In Sec. 5.2, we compare MicroDreamer with seven competitive baselines by utilizing the official code released in GitHub. See Tab. 2 for details about the URL and licenses of the assets we used in this paper.
Table 2: Codebase. URL and licenses are provided for the open-source assets we used in this paper.
Details in metric.
All the results presented in Sec. 5.2 were on the mesh representation. As TriplaneGaussian [33] has no official mesh export code, we apply the mesh exported code from LGM [30] for it. For extra refine time, we record the time of second stage refinement for MicroDreamer and DreamGaussian [9], the time of generating 3D from multi-view images for Wonder3D [39], and the time for meshing through a NeRF for LGM [30], respectively.
Data used in ablation study.
We chose seven images collected from the test dataset to perform our ablation study in Fig. 7 and Fig. 8. We will include these images in the submitted supplementary material.
Safeguards.
Our method, which requires no additional models and doesn’t release any new pre-trained model, can ensure security due to the inherent safe configuration of the employed 2D multi-view diffusion models [42, 43, 44].
Appendix E More results
Results on 3DGS.
We provide more results on mesh exported from 3DGS for MicroDreamer-20 iters in Fig. 12.
Figure 12: More results for MicroDreamer-20 iters.
Results on NeRF.
In Fig. 13, we present a comparison of the generated quality between SIR and SDS on NeRF in the same short time under different base diffusion models [42, 43, 6, 44]. SIR can generate NeRF of high visual quality in a shorter time than SDS, as the results represented. Notably, SIR accelerates the generation process in ImageDream more, as it’s an advanced multi-view diffusion model that can produce higher-quality multi-view images and support fewer discretization steps for the diffusion process, lowering NFEs.
Figure 13: Comparison of the generated quality between SIR and SDS in the same short time on NeRF. SIR can generate NeRF of high visual quality in less time than SDS.
Limitations.
In Fig. 14, we show some of the less successful cases generated by our method. Due to the limited ability of the diffusion model, inputs with complex geometric structures such as central hollowing may not result in satisfactory mesh outcomes. Furthermore, existing diffusion models struggle to generate the rear view of an object, which often leads to poorly textured back surfaces of the generated meshes. These problems may be solved as the quality of the images the multi-view diffusion model generated improves.
@@ Appendix E More results
Figure 14: Limitations. Visualization of some less satisfactory cases generated by our methods.
Appendix F Quantative comparison on 3DGS
We provide the detailed quantitative comparison table in Tab. 3. For a fair comparison, we also explored the performance of DreamGaussian with fewer iterations (300 iterations) whose generation time is near MicroDreamer-30 iters. As shown in Tab. 3, MicroDreamer performs better than DreamGaussian over a similar amount of time, further illustrating that MicroDreamer attains a commendable efficiency-performance balance.
Table 3: Quantitative comparisons. In terms of CLIP similarity, MicroDreamer achieves an excellent balance between efficiency and effectiveness.
Method CLIP-Similarity $\uparrow$ Generation time $\downarrow$ Extra refine $\downarrow$
Point-E 0.510 $\sim 24$s -
Shap-E 0.616 $\sim 5$s -
One-2-3-45 0.617 $\sim 42$s -
Wonder3D 0.699 $\sim 30$s $\sim 140$s
TriplaneGaussian 0.647 $\sim 7$s -
LGM 0.716 $\sim 3$s $\sim 50$s
DreamGaussian 0.698 $\sim 30$s $\sim 10$s
DreamGaussian-300 iters 0.656 $\sim 18$s $\sim 10$s
MicroDreamer-20 iters (Ours) 0.678 $\sim 18$s $\sim 4$s
MicroDreamer-30 iters (Ours) 0.687 $\sim 26$s $\sim 4$s

### v3

@@ VI-D Analysis and ablation study
(a) Results of single-step prediction using Eq. (2) for SDS.
(b) Results of applying SDS in pixel space.
@@
Figure 11: Failure of applying SDS in pixel space. SDS optimizing in pixel space generates poor results.
(a) Iterative reconstruction.
(b) Camera views.
(c) Time schedule.
(d) Forward process.
Figure 12: Detailed analyses. (a) Iterative reconstruction is necessary. (b) More camera views benefit. (c) A linear schedule is sufficient. (d) A hybrid forward process can be effective.
Analysis on optimization space. We implement SIR-latent on NeRF and compare its time consumption with SIR-pixel, as illustrated in Fig. 12a-b. We can see that SIR-pixel achieves 2-3 times more acceleration than SIR-latent, verifying the benefits of SIR in enabling optimization in pixel space. Contrary to SIR, we argue that SDS optimizing in pixel space is not effective even if it has the data predicting form by reparameterization as presented in Eq. (6). If we attempt to employ SDS for optimization in pixel space, i.e. optimizing 3D objects with 2D images like those generated by Eq. (2) in Fig. 11a, the resulting 3D content would be of subpar quality, as the results shown in Fig. 11b.
Ablation on iterative reconstruction. Fig. 12(a) shows the necessity of iterative reconstruction. The case of $K=0$ outputs our 3D initialization, which is reconstructed by a single process from images generated by random noise. When $K>30$, the generation quality doesn’t improve significantly. For efficiency, we set $K=20$ or $30$.
Ablation on number of camera poses. SIR necessitates more than 1 camera pose for reconstruction, as shown in Fig. 12(b), and the performance increases with more camera poses. We choose $6$ on 3DGS with Stable Zero123 [68] model to balance efficiency and quality. See Tab. I for values in other settings.
Ablation on time schedule. We compare our linear schedule, the random schedule employed in SDS [1], and a square schedule with $t_{2}^{(k)}=(\frac{k}{K})^{2}\cdot(0.9T-0.2T)+0.2T$. As shown in Fig. 12(c), the schedules perform similarly. For simplicity and efficiency, we use the linear schedule by default.
Ablation on forward process. Fig. 12(d) compares three different forward processes on 3DGS with the Stable Zero123 [68] model. The hybrid strategy performs better than noise-adding. As discussed in Sec. IV-C, we adopt the hybrid process rather than inversion for efficiency.
VI-E Limitation
We show some less successful cases generated by our method in Fig. 13. MicroDreamer faces challenges in generating complex geometric structures such as central hollows and may produce meshes with poor back surface textures. These problems may be mitigated as the quality of the generated multi-view images improves.
@@ VI-E Limitation
Figure 13: Limitation. Visualization of some less satisfactory cases generated by MicroDreamer.
VII Conclusion
We introduce SIR, an efficient and general algorithm combining the strengths of 3D reconstruction and iterative optimization to reduce total NFEs and enable optimization in pixel space in optimization-based 3D generation. SIR achieves a 5 to 20 times speed increase in NeRF generation compared to SDS. Remarkably, MicroDreamer generates high-quality meshes from 3DGS in about 20 seconds, outpacing the fastest optimization-based baseline DreamGaussian in quality and efficiency, and matching the speed of some feed-forward approaches with a competitive generation quality.
There is potential for further improving MicroDreamer’s efficiency via employing consistency models [88, 89] or alternative sampling models that require fewer steps [90, 91]. Additionally, the fidelity and 3D consistency of the objects produced by MicroDreamer are directly limited by the quality of the outputs from multi-view diffusion. Nevertheless, we believe SIR is promising and may inspire future work as the multi-view diffusion evolves.
Furthermore, MicroDreamer can provide artists with the convenience of creating 3D assets. However, as a generative approach, our method may also be used for fabricating data and news. Moreover, the 3D content generated from input images may infringe on the privacy or copyright of others. While automatic detection might mitigate these issues, they warrant careful consideration.